OpenStudy (anonymous):

7y^2-y-1/2=0 y=? smaller value y=? larger value

OpenStudy (anonymous):

\[7y^2 - y - (1/2) = 0\] \[14y^2 - 2y - 1 = 0\] \[y= \left[2 \pm \sqrt{4 - 4(14)(-1)}\right]\left(\frac{1}{28}\right)\] \[y= \left[2 \pm \sqrt{60}\right]\left(\frac{1}{28}\right)\] \[y= \left[2 \pm 2\sqrt{15}\right]\left(\frac{1}{28}\right)\] \[y= \left[1 \pm \sqrt{15}\right]\left(\frac{1}{14}\right)\] so \[y = \frac{1}{14} + \frac{\sqrt{15}}{14}\] (the larger one) and \[y = \frac{1}{14} - \frac{\sqrt{15}}{14}\] (the smaller one)